OpenStudy (anonymous):

A right triangle has one side of length 7 and a hypotenuse of length 10. How long is the third side of the triangle?

OpenStudy (anonymous):

a^2+b^2=c^2 pythagorean's theorem, plug and chug

OpenStudy (anonymous):

can you do the rest?

OpenStudy (anonymous):

Oh I get it! I think

OpenStudy (anonymous):

So like 7^2+10^2 = c^2?

OpenStudy (anonymous):

alright, the equation is 10^2=7^2+b^2 incase you need to check your work 100 = 49 +b^2 b^2 = 51 b = sqrt(51)

OpenStudy (anonymous):

c has to be hypotenuse, a and b are the adjacent/opposite sides of the right triangle

OpenStudy (anonymous):

Oh and all I have to do is square root 51?

OpenStudy (anonymous):

yep, do you see how i arrived at it though?

OpenStudy (anonymous):

Yeah I think so. Since a and b are adjacent sides and c is the hypotenuse you just plug in what you already have which is a and c

OpenStudy (anonymous):

you got it

OpenStudy (anonymous):

and the rest of it works out

OpenStudy (anonymous):

Thank you! :D